# Multiple Information Prompt Learning for Cloth-Changing Person Re-Identification

### v1

@@ Abstract
Abstract - Cloth-changing person re-identification is a subject closer to the real world, which focuses on solving the problem of person re-identification after pedestrians change clothes. The primary challenge in this field is to overcome the complex interplay between intra-class and inter-class variations and to identify features that remain unaffected by changes in appearance. Sufficient data collection for model training would significantly aid in addressing this problem. However, it is challenging to gather diverse datasets in practice. Current methods focus on implicitly learning identity information from the original image or introducing additional auxiliary models, which are largely limited by the quality of the image and the performance of the additional model. To address these issues, inspired by prompt learning, we propose a novel multiple information prompt learning (MIPL) scheme for cloth-changing person ReID, which learns identity robust features through the common prompt guidance of multiple messages. Specifically, the clothing information stripping (CIS) module is designed to decouple the clothing information from the original RGB image features to counteract the influence of clothing appearance. The Bio-guided attention (BGA) module is proposed to increase the learning intensity of the model for key information. A dual-length hybrid patch (DHP) module is employed to make the features have diverse coverage to minimize the impact of feature bias. Extensive experiments demonstrate that the proposed method outperforms all state-of-the-art methods on the LTCC, Celeb-reID, Celeb-reID-light, and CSCC datasets, achieving rank-1 scores of 74.8%, 73.3%, 66.0%, and 88.1%, respectively. When compared to AIM (CVPR23), ACID (TIP23), and SCNet (MM23), MIPL achieves rank-1 improvements of 11.3%, 13.8%, and 7.9%, respectively, on the PRCC dataset.  [^1] Manuscript received June-27th, 2024; This work was supported in part by the National Natural Science Foundation of China (No.62372325, No.61872270), Natural Science Foundation of Tianjin Municipality (No.23JCZDJC00280), Shandong project towards the integration of education and industry (No.2022JBZ01-03). S.X Wei, Z. Gao and C.J Ma are with the Shandong Artificial Intelligence Institute, Qilu University of Technology (Shandong Academy of Sciences), Jinan, 250014, P.R China. S.X Wei, Z. Gao and Y.B Zhao are with Key Laboratory of Computer Vision and System, Ministry of Education, Tianjin University of Technology, Tianjin, 300384, P.R China. W.L Guan is with the School of Electrontics and Information Engineering, Harbin Institute of Technology, Shenzhen, 15001, P.R China. .
@@ I Introduction
Person Re-identification (ReID) is a prominent research focus within the fields of computer vision and machine learning. It aims to solve the problem of retrieving target pedestrians across non-overlapping cameras, and determine whether the images taken by different cameras contain a specific pedestrian. This task plays an important role in areas such as public safety, smart commerce, and smart security (e.g., identifying customers for personalization, tracking potential criminal suspects). Traditional person ReID methods [28, 29, 44, 37, 5, 13, 20] usually rely on pedestrians wearing the same clothes in different shots. However, in real-world scenarios, pedestrian appearance can change significantly over time, making it difficult for models to learn effective identification features. This variability presents a substantial challenge for person ReID. To address this issue, researchers are increasingly focusing on the cloth-changing person re-identification (CC-ReID) task. This approach aims to identify other identity-related features that are independent of clothing, such as body posture and gait. These features can still effectively identify pedestrians even when their clothing changes.
Some researchers [1, 8, 35, 9, 30, 7] have made active attempts in person ReID task under changing clothes scenario. The public datasets PRCC [34], LTCC [24], Celeb-reID [16], Celeb-reid-light [15] and CSCC [33] for person ReID research have also been constructed. Figure 1 shows some examples of CC-ReID images, where the first row represents the same person wearing different clothes and the second row represents different people wearing similar clothes, from which we can observe that the first row exhibits a large difference in pedestrian appearance and the second row shows a very small difference in pedestrian appearance. Under CC-ReID, the appearance of human clothing can no longer be used as the main basis for recognition but has become an interference factor. At this time, the traditional ReID method that largely depends on the appearance characteristics of clothing is no longer competent, so the ReID method for changing clothes is a more challenging and urgent problem to be solved. In CC-ReID, because of the unreliability of clothing information, the intuitive response is to model identity information unrelated to clothing. Recently, some novel methods for changing person ReID have been proposed, for example, Jin et al. [18] proposed a Gait-assisted Image-based cloth-changing ReID (GI-ReID) framework, which introduces Gait information to assist in learning clothing independent representations. Gao et al. [7] proposed a novel multigranular visual-semantic embedding algorithm (MVSE) to deeply explore the visual semantic information and pedestrian attribute information, so as to effectively solve the problem of CC-ReID.
@@
To address the above issues, we design a new MIPL framework to learn identity robust features using the prompt guidance of multiple information. For problem (1), inspired by [20], we introduce a phased vision-language learning strategy to effectively establish the correspondence between visual representation and high-level language description, and design a clothing information stripping module to effectively decouple clothing information from original image features with the prompt of clothing attribute text description. For problem (2), the biological information guidance branch is proposed, which explicitly prompts the model to learn biological key features with strong identity correlation through local unique biological information attention prompts. In addition, for problem (3), a dual-length hybrid patch module is designed to reduce the impact of feature deviation and make the features have diverse coverage. The main contributions of this paper are summarized as follows.
@@ II Related Work
II-A Clothing-Consistent Person ReID
The datasets Market-1501 [43], MSMT17 [31], CUHK03 [21] used by most person ReID methods are collected in a short period of time, and they are all carried out under the assumption that the appearance of pedestrians’ clothes will not change. They mainly focus on traditional factors such as illumination, pose, viewpoint, and occlusion, and achieve satisfactory performance. For example, relying on the powerful modeling ability of CNN network [12], Zhou et al. [44] designed a new omni-scale network (OSNet) to capture features at different spatial scales, and to collaboratively combine multi-scale features to obtain full-scale pedestrian representation. Ye et al. [37] conducted a comprehensive review and in-depth analysis of the existing person ReID technologies from the perspectives of deep feature representation learning, deep metric learning and ranking optimization. In addition, a baseline method of Attention Generalized mean pooling with Weighted triplet loss (AGW) is proposed. Gao et al. [5] proposed a deep spatial pyramid feature collaborative reconstruction model (DCR), which uses joint reconstruction to solve the problems of pedestrian pose, perspective changes and occlusion in ReID. With the success of Transformers [4] in vision tasks, He et al. [13] propose a pure transformer-based baseline framework (TransRelD), which encodes edge information such as viewpoint and camera through learnable embeddings, and rearranges patch embeddings to generate features with better discriminative power. Recently, vision-language models [25] have emerged and achieved success in many downstream tasks, Li et al. [20] first introduced it into ReID task and proposed Clip-ReID, which uses a learnable prompt text to align image features to optimize the distance between objects. These methods can effectively deal with traditional factors such as illumination, pose, perspective and occlusion, but are limited to completing the person ReID task in a short period of time (that is, under the assumption that the person’s clothing appearance does not change).
@@ II-B Cloth-changing Person ReID
In the real scenario, the assumption that the appearance of clothes does not change over time is quickly broken, and the clothing-consistent person ReID method is no longer efficient. Therefore, researchers begin to establish some cloth-changing person ReID datasets [38, 34, 24, 15, 16], and pay attention to the task of cloth-changing person ReID. For example, Yu et al. [38] constructed a novel change of clothing recognition benchmark COCAS, which uses clothing templates and pedestrian images for combination to search the target image. Chen et al. [1] combined person ReID and 3D human reconstruction to propose an end-to-end architecture for 3D shape learning (3DSL) to extract texture-insensitive 3D shape embeddings directly from 2D images. Hone et al. [14] proposed a two-stream Fine-grained shape-appearance Mutual learning (FSAM) framework to supplement clothing-independent knowledge from shape streams to appearance features through dense interactive mutual learning. Chen et al. [2] proposed a multi-scale appearance and contour deep infomax(MAC-DIM) module to maximize the mutual information between color appearance features and wheel shape features to overcome the problem of clothing color bias. Gu et al. [8] designed a cloths-based Adversarial Loss (CAL) to penalize the model’s ability to predict Clothes, mining features unrelated to clothes from the original RGB images. Zhao et al. [42] regard clothing change as a fine-grained domain/style transfer, and propose a joint Identity-aware Mixstyle and Graph-enhanced Prototype for cloth-changing person Re-ID. Liu et al. [23] proposed a Pose-Guided Attention Learning (PGAL) framework, which uses a pose estimation network to align keypoint features and improves the feature representation of non-keypoint regions of the human body through multi-head self-attention. Gao et al. [6] proposed a new semantic-aware attention and visual shielding network (SAVS), which shields the cues related to clothing appearance and only focuses on visual semantic information that is insensitive to viewpoint/pose changes. Yang et al. [35] designed a causality-based Auto Intervention Model (AIM) for cloth-changing person ReID, which captures clothing bias and identity cues separately, and simulates causal intervention by stripping clothing inference from identity representation learning. Guo et al. [9] proposed a Semantic-aware Consistency Network (SCNet) to learn identity-related Semantic features by proposing effective consistency constraints. Wang et al. [30] focus on continuous shape distribution at pixel level and propose Continuous Surface Correspondence Learning (CSCL) to enhance the global understanding of human body shape by shape embedding paradigm based on 2D-3D correspondence. Cui et al. [3] proposed Deep Component Reconstruction Re-ID (DCR-ReID), which realizes the controllable decoupling of clothes-independent features and clothes-related features. Yang et al. [36] focus on potential identity cues hidden in appearance and structural features, and propose an Auxiliary free Competitive IDentification (ACID) model to perform precise identity identification without auxiliary data. Wu et al. [32] proposed a two-stream hybrid Convolution Transformer Network (CT-Net), which combines CNN and Transformer in parallel in an end-to-end learning scheme. Han et al. [10] aim to scale up the training data and propose a novel clothing-change Feature Augmentation (CCFA) model that implicitly integrates semantically meaningful Clothing Change augmentation in the feature space. Liu et al. [22] proposed a Dual-Level Adaptive Weighting (DLAW) solution to measure the degree of cloth-changing and evaluate the influence of image and feature levels on cloth-changing patterns, thereby addressing the cloth-changing ReID problem. Zhang et al. [39] proposed the Multi-Biometric Unified Network (MBUNet) framework, which leverages multi-biometric features to learn cloth-changing cues that are unrelated to clothing. Additionally, mAP optimization is introduced to address the discrete optimization problem of mAP. To alleviate the dependence on rich data collection, Jia et al. [17] enhanced the feature learning process by designing a powerful complementary data augmentation strategy called Pos-Neg. This strategy collaboratively promotes the model to learn more robust and discriminative representations without the need for additional information or the cost of expanding the latent sample space. Due to the large amount of redundant spatial information in visual images [11], the existing methods cannot decouple this information well, resulting in poor performance of the existing methods. Therefore, in this work, we will focus on several aspects (i.e., i. Control the spatial redundant information, ii. Use more identity cues, iii. reducing feature bias) energetically explore robust representations of a person with different clothing.
@@ III Multiple Information Prompt Learning Network
Figure 2: Pipeline of the proposed MIPL approach. It consists of the backbone, the CIS module, the BGA module, and the DHP module. ’SCHP’ is a semantic analysis module to obtain the human semantic information. ’Norm.’ and ’Trans.’ denote the normalization and transpose operations, respectively.
In this work, we develop a new MIPL algorithm for cloth-changing person ReID, which applies multiple aspects of information to co-prompts guided model learning. It can decouple redundant information in visual modes, make full use of easily captured pedestrian identity clues, reduce the influence of feature bias, and enable the model to learn the features of identity robustness. MIPL network is mainly composed of CIS module, BGA module and DHP module. The framework is shown in Figure 2. Specifically, the model is divided into two stages. In the first stage, only the CIS module and backbone network participate in the training, freezing the parameters of the image and the text encoder, optimizing a set of learnable text prompt words for each identity and clothing; In the second stage, the BGA module and DHP module are added to freeze the text encoder and optimize the text prompt words, and fine-tune the image encoder. In addition, the clothing image of the CIS module and the biological information image of the BGA module are both obtained with the assistance of the human parsing model SCHP [19], and the output of the backbone network is fed to the DHP module to learn the features of diverse coverage. It is important that they learn together in a unified framework. In the following sections, we will cover each module and the loss function separately.
@@ III-A Clothing Information Stripping (CIS)
In order to effectively decouple the visual redundant information in the visual modality, inspired by the two-stage prompt learning [25], we pre-train the learnable text prompt words of identity and clothing to supplement the text information, to establish an effective correspondence between visual representations and high-level language descriptions, and constrain the model to accurately locate the clothing area through the text description to decouple the clothing area from the non-clothing area. It reduces the influence of clothing information on CC-ReID task. Specifically, in the first training stage, A set of learnable prompt words are introduced, which are an identity-dependent text prompt (”A photo of a $[X]_{1}^{p}$ $[X]_{2}^{p}$... $[X]_{M}^{p}$ person.”) and a clothes-dependent text prompt (”A photo of a $[X]_{1}^{c}$ $[X]_{2}^{c}$... $[X]_{M}^{c}$ clothes.”) Where each $[x]_{*}^{*}$is a learnable text token with the same dimension as the embedded word,$p$ denotes the text token belonging to the identity dependency, $c$ denotes the text token belonging to the clothing dependency, and $M$ denotes the number of learnable text tokens. Then we use the text encoder and image encoder with frozen parameters to obtain the corresponding text features $F_{ori}^{text}$, $F_{clo}^{text}$ and image features $F_{ori}^{img}$, $F_{clo}^{img}$ (the encoder is pre-trained by CLIP [25]). A contrastive learning loss function is used to constrain the alignment between text features and image features. In this way, a unique prompt is learned for different identities and clothes separately, providing precise guidance for the decoupling of clothing information from the original image. In the second training stage, we only optimize the image encoder and freeze the trained text prompt words and the parameters of the text encoder. The trained text features are used to align the clothing and body regions, and then the clothing stripping loss is designed to decouple the clothing information from the identity information in the image. In addition, the spatial consistency loss is introduced to regularize the clothing mapping features (obtained by passing the original features through the last block of the image encoder) and clothing features into a unified feature hypersphere, so that the clothing features can be better aligned, to complete the feature stripping. It should be noted that the above feature operation details will be given in the loss function.
III-B Bio-guided Attention (BGA)
where $\mathcal{N}$ represents the normalization operation, $T$ represents the transpose operation, $\otimes$ represents matrix multiplication, and $\mathcal{M}$ indicates the biological key information mask. In order to effectively use the information enhancement features, the biological guided loss is used to transfer the knowledge of biological key area information to the backbone network branches, which prompts the model to strengthen the learning of strong identity-related regions. In this way, effective prompting models are biologically information-oriented to learn biologically key features that are strongly correlated with identity.
III-C Dual-length Hybrid Patch (DHP)
III-D Loss Function
$\mathcal{L}_{Stage1}=\mathcal{L}_{i2t}^{ID}(y_{i})+\mathcal{L}_{i2t}^{C}(y_{c})+\mathcal{L}_{t2i}^{ID}(y_{i})+\mathcal{L}_{t2i}^{C}(y_{c}),$ (3)
which includes the image-text contrastive loss $\mathcal{L}_{{i2t}}$ and the text-image contrastive loss $\mathcal{L}_{{t2i}}$, where $ID$ and $C$ represent identity and clothing dependency respectively, and $y_{i}$ represents identity labels, and $y_{c}$ represents clothing labels. For the image-text contrastive loss is calculated as:
@@ III-D Loss Function
Among them, $P(y_{i})$ is the set of indices of all positives for $T_{y_{i}}$ in the batch, and ||$\cdot$|| ⋅ | is its cardinality. In addition, for $\mathcal{L}_{i2t}^{C}(y_{c})$ and $\mathcal{L}_{t2i}^{C}(y_{c})$ follow the above calculation.
The second training stage. In this stage, both the text prompt and the text encoder are frozen and only the image encoder is optimized, we follow previous ReID work [13, 20, 44] to calculate cross-entropy loss and triplet loss to optimize the image encoder. In addition, in order to strip the clothing interference information, we design the clothing stripping loss to guide the model to decouple and strip the clothing information. In order to effectively use biological information to enhance features, the biological guided loss is used to prompt the model to learn biological knowledge with a strong identity correlation. Thus, the loss function in the second stage is defined by
For the cross-entropy loss $\mathcal{L}_{ce}$ and the triplet loss $\mathcal{L}_{tri}$, they are calculated as follows:
$\mathcal{L}_{ce}=\frac{1}{B}\sum_{i=1}^{B}-\log p(x\mid y)$ (7)
$\mathcal{L}_{tri}=\frac{1}{B}\sum_{i=1}^{B}\max\{m+d_{p}-d_{n},0\}$ (8)
Where $p(x\mid y)$ is the predicted probability that the sample $x$ belongs to the ground truth $y$. $d_{p}$ and $d_{n}$ are the feature distances of positive pair and negative pair, and $m$ is the margin of $\mathcal{L}_{tri}$. For the clothing stripping loss $\mathcal{L}_{CS}$, the process is defined as follows:
$\mathcal{L}_{CS}=\mathcal{L}_{Guide}+\mathcal{L}_{sc}+\mathcal{L}_{de},$ (9)
Specifically, in the CIS model, the image-to-text cross-entropy loss is calculated using the text embeddings trained in the first stage (including $N_{i}$ identity text embeddings and $N_{c}$ clothing text embeddings), and the image features are aligned by the text embeddings to guide the model to extract more accurate features, and the guide loss is calculated as:
@@
In order to make more comprehensive use of identity information, the biological information enhanced features are used to prompt the model to learn the biological knowledge with a strong identity, and the biological knowledge with a strong identity is transferred to the original features. The biological guided loss $\mathcal{L}_{BG}$ is calculated as:

### v2

@@ Abstract
Abstract - Cloth-changing person re-identification is a subject closer to the real world, which focuses on solving the problem of person re-identification after pedestrians change clothes. The primary challenge in this field is to overcome the complex interplay between intra-class and inter-class variations and to identify features that remain unaffected by changes in appearance. Sufficient data collection for model training would significantly aid in addressing this problem. However, it is challenging to gather diverse datasets in practice. Current methods focus on implicitly learning identity information from the original image or introducing additional auxiliary models, which are largely limited by the quality of the image and the performance of the additional model. To address these issues, inspired by prompt learning, we propose a novel multiple information prompt learning (MIPL) scheme for cloth-changing person ReID, which learns identity robust features through the common prompt guidance of multiple messages. Specifically, the clothing information stripping (CIS) module is designed to decouple the clothing information from the original RGB image features to counteract the influence of clothing appearance. The bio-guided attention (BGA) module is proposed to increase the learning intensity of the model for key information. A dual-length hybrid patch (DHP) module is employed to make the features have diverse coverage to minimize the impact of feature bias. Extensive experiments demonstrate that the proposed method outperforms all state-of-the-art methods on the LTCC, Celeb-reID, Celeb-reID-light, and CSCC datasets, achieving rank-1 scores of 74.8%, 73.3%, 66.0%, and 88.1%, respectively. When compared to AIM (CVPR23), ACID (TIP23), and SCNet (MM23), MIPL achieves rank-1 improvements of 11.3%, 13.8%, and 7.9%, respectively, on the PRCC dataset.  [^1] Manuscript received June-27th, 2024; This work was supported in part by the National Natural Science Foundation of China (No.62372325, No.62402255), Natural Science Foundation of Tianjin Municipality (No.23JCZDJC00280), Shandong Provincial Natural Science Foundation (No.ZR2024QF020), Shandong Province National Talents Supporting Program (No.2023GJJLJRC-070), Shandong project towards the integration of education and industry (No.801822020100000024), Young Talent of Lifting engineering for Science and Technology in Shandong (No. SDAST2024QTB001), Shandong Project towards the Integration of Education and Industry (No.2024ZDZX11). S.X Wei, Z. Gao, Y.B Zhao and S.Y Chen are with Key Laboratory of Computer Vision and System, Ministry of Education, Tianjin University of Technology, Tianjin, 300384, P.R China. C.J Ma and Z. Gao are with the Shandong Artificial Intelligence Institute, Qilu University of Technology (Shandong Academy of Sciences), Jinan, 250014, P.R China. W.L Guan is with the School of Electronics and Information Engineering, Harbin Institute of Technology, Shenzhen, 150001, P.R China.
@@ I Introduction
Person Re-identification (ReID) is a prominent research focus within the fields of computer vision and machine learning. It aims to solve the problem of retrieving target pedestrians across non-overlapping cameras, and determine whether the images taken by different cameras contain a specific pedestrian. This task plays an important role in areas such as public safety, smart commerce, and smart security (e.g., identifying customers for personalization, tracking potential criminal suspects). Traditional person ReID methods [32, 33, 26, 29, 7, 15, 49, 42, 22] usually rely on pedestrians wearing the same clothes in different shots. However, in real-world scenarios, pedestrian appearance can change significantly over time, making it difficult for models to learn effective identification features. This variability presents a substantial challenge for person ReID. To address this issue, researchers are increasingly focusing on the cloth-changing person re-identification (CC-ReID) task. This approach aims to identify other identity-related features that are independent of clothing, such as body posture and gait. These features can still effectively identify pedestrians even when their clothing changes.
Some researchers [2, 10, 9, 1, 40, 11, 34] have made active attempts in person ReID task under changing clothes scenario. The public datasets PRCC [38], LTCC [27], Celeb-reID [18], Celeb-reid-light [17] and CSCC [37] for person ReID research have also been constructed. Figure 1 shows some examples of CC-ReID images, where the first row represents the same person wearing different clothes and the second row represents different people wearing similar clothes, from which we can observe that the first row exhibits a large difference in pedestrian appearance and the second row shows a very small difference in pedestrian appearance. Under CC-ReID, the appearance of human clothing can no longer be used as the main basis for recognition but has become an interference factor. At this time, the traditional ReID method that largely depends on the appearance characteristics of clothing is no longer competent, so the ReID method for changing clothes is a more challenging and urgent problem to be solved. In CC-ReID, because of the unreliability of clothing information, the intuitive response is to model identity information unrelated to clothing. Recently, some novel methods for changing person ReID have been proposed, for example, Jin et al. [20] proposed a Gait-assisted Image-based cloth-changing ReID (GI-ReID) framework, which introduces Gait information to assist in learning clothing independent representations. Gao et al. [9] proposed a novel multigranular visual-semantic embedding algorithm (MVSE) to deeply explore the visual semantic information and pedestrian attribute information, so as to effectively solve the problem of CC-ReID.
@@
To address the above issues, we design a new MIPL framework to learn identity robust features using the prompt guidance of multiple information. For problem (1), inspired by [22], we introduce a multi-stage vision-language learning strategy to effectively establish the correspondence between visual representation and high-level language description, and design a clothing information stripping module to effectively decouple clothing information from original image features with the prompt of clothing attribute text description. For problem (2), the biological information guidance branch is proposed, which explicitly prompts the model to learn biological key features with strong identity correlation through local unique biological information attention prompts. In addition, for problem (3), a dual-length hybrid patch module is designed to reduce the impact of feature deviation and make the features have diverse coverage. The main contributions of this paper are summarized as follows.
@@ II Related Work
II-A Cloth-consistent Person ReID
In the field of person ReID, the characteristics of commonly used datasets and the limitations of related methods are worthy of in-depth discussion. The datasets Market-1501 [48], MSMT17 [35], and CUHK03 [23] used by most person ReID methods are collected in a short period of time, and they are all carried out under the assumption that the appearance of pedestrians’ clothes will not change. These datasets mainly focus on traditional factors such as illumination, pose, viewpoint, and occlusion and achieve satisfactory performance. For instance, relying on the powerful modeling ability of the CNN network [14], Zhou et al. [49] designed a new omni-scale network (OSNet). This network emphasizes multi-scale feature capture and combination to achieve better representation. However, OSNet may have higher computational complexity due to its multi-scale feature capture design, and there is a risk of overfitting with its intricate architecture. Meanwhile, Ye et al. [42] offers a comprehensive analysis and new baseline method in person reID. Reviews and analyzes existing person ReID techniques from deep feature representation learning, deep metric learning, and ranking optimization. Gao et al. [7] proposed a deep spatial pyramid feature collaborative reconstruction model (DCR), which uses joint reconstruction to address the issues of pedestrian pose, perspective changes, and occlusion in ReID. It provides effective support for person ReID tasks in scenarios with these challenges. With the success of Transformers [6] in vision tasks, He et al. [15] proposed a pure transformer-based baseline framework (TransReID). This framework encodes edge information such as viewpoint and camera through learnable embeddings and rearranges patch embeddings to generate more discriminative features.
These methods can effectively deal with traditional factors such as illumination, pose, perspective and occlusion, but are limited to completing the person ReID task in a short period of time (that is, under the assumption that the person’s clothing appearance does not change).
@@ II-B Cloth-changing Person ReID
In the real scenario, the assumption that the appearance of clothes does not change over time is quickly broken, and the cloth-consistent person ReID method is no longer efficient. Therefore, researchers begin to establish some cloth-changing person ReID datasets [17, 43, 27, 18, 38, 39], and pay attention to the task of cloth-changing person ReID. For example, Yu et al. [43] constructed a novel change of clothing recognition benchmark COCAS, which uses clothing templates and pedestrian images for combination to search the target image. Additionally, in order to reduce the reliance on extensive data collection, Jia et al. [19] enhanced the feature learning process by devising a potent complementary data augmentation strategy named Pos-Neg. This strategy jointly prompts the model to learn more robust and discriminative representations without requiring additional information or incurring the cost of expanding the latent sample space. Han et al. [12] aim to enlarge the training data and put forward a novel clothing-change Feature Augmentation (CCFA) model that implicitly integrates semantically meaningful Clothing Change augmentation in the feature space. Liu et al. [24] proposed a Dual-Level Adaptive Weighting (DLAW) solution to measure the degree of cloth-changing and assess the influence of image and feature levels on cloth-changing patterns, thereby resolving the cloth-changing ReID problem. Merely relying on data augmentation cannot comprehensively and effectively express the targets in images. Therefore, researchers have become keen on using auxiliary information to enable the model to learn jointly. So, Chen et al. [2] combined person ReID and 3D human reconstruction to propose an end-to-end architecture for 3D shape learning (3DSL) to extract texture-insensitive 3D shape embeddings directly from 2D images. Subsequently, Hone et al. [16] proposed a two-stream Fine-grained shape-appearance Mutual learning (FSAM) framework to supplement clothing-independent knowledge from shape streams to appearance features through dense interactive mutual learning. Moreover, Chen et al. [3] proposed a multi-scale appearance and contour deep infomax(MAC-DIM) module to maximize the mutual information between color appearance features and wheel shape features to overcome the problem of clothing color bias. Liu et al. [25] proposed a Pose-Guided Attention Learning (PGAL) framework, which uses a pose estimation network to align keypoint features and improves the feature representation of non-keypoint regions of the human body through multi-head self-attention. Zhang et al. [44] proposed the Multi-Biometric Unified Network (MBUNet) framework, which leverages multi-biometric features to learn cloth-changing cues that are unrelated to clothing. Wang et al. [34] focus on continuous shape distribution at pixel level and propose Continuous Surface Correspondence Learning (CSCL) to enhance the global understanding of human body shape by shape embedding paradigm based on 2D-3D correspondence. These methods all use additional modal information (3D, shape, and contour, etc) to assist the model in learning. However, this inevitably introduces uncontrollable overhead and noise. As a result, a batch of methods that deeply explore the implicit information contained within the image itself have emerged. For example, Gu et al. [10] designed a clothes-based Adversarial Loss (CAL) to penalize the model’s ability to predict clothes, mining features unrelated to clothes from the original RGB images. Zhao et al. [47] regard clothing change as a fine-grained domain/style transfer, and propose a joint Identity-aware Mixstyle and Graph-enhanced Prototype for cloth-changing person Re-ID. Gao et al. [8] proposed a new semantic-aware attention and visual shielding network (SAVS), which shields the cues related to clothing appearance and only focuses on visual semantic information that is insensitive to viewpoint/pose changes. Yang et al. [40] designed a causality-based Auto Intervention Model (AIM) for cloth-changing person ReID, which captures clothing bias and identity cues separately, and simulates causal intervention by stripping clothing inference from identity representation learning. Guo et al. [11] proposed a Semantic-aware Consistency Network (SCNet) to learn identity-related semantic features by proposing effective consistency constraints. Cui et al. [4] proposed Deep Component Reconstruction Re-ID (DCR-ReID), which realizes the controllable decoupling of clothes-independent features and clothes-related features. Yang et al. [41] focus on potential identity cues hidden in appearance and structural features, and propose an Auxiliary-free Competitive IDentification (ACID) model to perform precise identity identification without auxiliary data. Additionally, Wu et al. [36] proposed a two-stream hybrid Convolution Transformer Network (CT-Net), which combines CNN and Transformer in parallel in an end-to-end learning scheme.
Due to the large amount of redundant spatial information in visual images [13], the existing methods cannot decouple this information well, resulting in poor performance of the existing methods. Therefore, in this work, we will focus on several aspects (i.e., i. Control the spatial redundant information, ii. Use more identity cues, iii. reducing feature bias) energetically explore robust representations of a person with different clothing.
@@ III Multiple Information Prompt Learning Network
Figure 2: Pipeline of the proposed MIPL approach. It consists of the backbone, the CIS module, the BGA module, and the DHP module. It is a two-stage model with stage 1 and stage 2. In stage 1, the text prompts (Clothing prompt and Identity prompt) are optimized, and in stage 2, the image encoder is trained. ’SCHP’ is a semantic analysis module to obtain the human semantic information. ’Norm.’ and ’Trans.’ denote the normalization and transpose operations, respectively. Other marks are marked in the lower right corner of pipeline.
In this work, we develop a new MIPL algorithm for cloth-changing person ReID, which applies multiple aspects of information to co-prompts guided model learning. It can decouple redundant information in visual modes, make full use of easily captured pedestrian identity clues, reduce the influence of feature bias, and enable the model to learn the features of identity robustness. MIPL network is mainly composed of CIS module, BGA module and DHP module. The framework is shown in Figure 2. Specifically, the model is divided into two stages. In the first stage, only the CIS module and backbone network participate in the training, freezing the parameters of the image and the text encoder, optimizing a set of learnable text prompt words for each identity and clothing; In the second stage, the BGA module and DHP module are added to freeze the text encoder and optimized the text prompt words, and fine-tune the image encoder. In addition, the clothing image of the CIS module and the biological information image of the BGA module are both obtained with the assistance of the human parsing model SCHP [21], and the output of the backbone network is fed to the DHP module to learn the features of diverse coverage. It is important that they learn together in a unified framework. In the following sections, we will cover each module and the loss function separately.
@@ III-A Clothing Information Stripping (CIS)
In order to effectively decouple the visual redundant information in the visual modality, inspired by the two-stage prompt learning [28], in MIPL, the image encoder is implemented as a Transformer architecture like Vit-B/16 [6] to generate image representations. The text encoder is implemented as a language Transformer architecture such as BERT [5] to generate text representations. Next, the image features and text features are normalized and linearly projected into the cross-modal embedding space for visual language contrastive learning. To be specific, we pre-train the learnable text prompt words of identity and clothing to supplement the text information, to establish an effective correspondence between visual representations and high-level language descriptions, and constrain the model to accurately locate the clothing area through the text description to decouple the clothing area from the non-clothing area. It reduces the influence of clothing information on CC-ReID task. Specifically, in the first training stage, A set of learnable prompt words are introduced, which are an identity-dependent text prompt (”A photo of a $[X]_{1}^{p}$ $[X]_{2}^{p}$... $[X]_{M}^{p}$ person.”) and a clothes-dependent text prompt (”A photo of a $[X]_{1}^{c}$ $[X]_{2}^{c}$... $[X]_{M}^{c}$ clothes.”) Where each $[x]_{*}^{*}$is a learnable text token with the same dimension as the embedded word,$p$ denotes the text token belonging to the identity dependency, $c$ denotes the text token belonging to the clothing dependency, and $M$ denotes the number of learnable text tokens. Then we use the text encoder and image encoder with frozen parameters to obtain the corresponding text features $F_{ori}^{text}$, $F_{clo}^{text}$ and image features $F_{ori}^{img}$, $F_{clo}^{img}$ (the encoder is pre-trained by CLIP [28]). A contrastive learning loss function is used to constrain the alignment between text features and image features.
For the image-text contrastive loss is calculated as:
@@
Among them, $P(y_{i})$ is the set of indices of all positives for $T_{y_{i}}$ in the batch, and ||$\cdot$|| ⋅ | is its cardinality. In addition, for $\mathcal{L}_{i2t}^{C}(y_{c})$ and $\mathcal{L}_{t2i}^{C}(y_{c})$ follow the above calculation. In this way, a unique prompt is learned for different identities and clothes separately, providing precise guidance for the decoupling of clothing information from the original image.
In the second training stage, we only optimize the image encoder and freeze the trained text prompt words and the parameters of the text encoder. The trained text features are used to align the clothing and body regions, and then the clothing stripping loss is designed to decouple the clothing information from the identity information in the image.
Specifically, the image-to-text cross-entropy loss is calculated using the text embeddings trained in the first stage (including $N_{i}$ identity text embeddings and $N_{c}$ clothing text embeddings), and the image features are aligned by the text embeddings to guide the model to extract more accurate features, and the guide loss is calculated as:
@@
For the clothing stripping loss $\mathcal{L}_{CS}$, the process is defined as follows:
$\mathcal{L}_{CS}=\mathcal{L}_{Guide}+\mathcal{L}_{sc}+\mathcal{L}_{de}$ (6)
In addition, the spatial consistency loss is introduced to regularize the clothing mapping features (obtained by passing the original features through the last block of the image encoder) and clothing features into a unified feature hypersphere, so that the clothing features can be better aligned, to complete the feature stripping. It should be noted that the above feature operation details will be given in the loss function.
III-B Bio-guided Attention (BGA)
where $\mathcal{N}$ represents the normalization operation, $T$ represents the transpose operation, $\otimes$ represents matrix multiplication, and $\mathcal{M}$ indicates the biological key information mask. In order to effectively use the information enhancement features, the biological guided loss is used to transfer the knowledge of biological key area information to the backbone network branches, which prompts the model to strengthen the learning of strong identity-related regions. The biological guided loss $\mathcal{L}_{BG}$ is calculated as: